OpenStudy (anonymous):

solve tan3x - sec3x = 5

OpenStudy (jhannybean):

rearange so you have \[\ \tan(3x) -5 = \sec (3x)\]\[\ ( \tan(3x) -5)^2 = (\sec(3x))^2\]\[\ \tan^2 (3x) -10\tan(3x) - 25 =\sec^2 (3x)\] How far I've gotten.

OpenStudy (jhannybean):

Im stuck here :|

OpenStudy (mathmath333):

\(\Large sec^23x-tan^23x=1\)

OpenStudy (jhannybean):

How did you get that? :o

OpenStudy (mathmath333):

\(\Large sec^2A-tan^2A=1\) famous identity http://en.wikipedia.org/wiki/List_of_trigonometric_identities

OpenStudy (jhannybean):

Oh this, yes.

OpenStudy (jhannybean):

I still don't understand how to apply it :(

OpenStudy (anonymous):

Try factoring a \(\sec x\).

OpenStudy (anonymous):

\[\begin{align*} \tan3x-\sec3x&=5\\ \frac{\sin3x}{\cos3x}-\frac{1}{\cos3x}&=5\\ \frac{1}{\cos3x}\left(\sin3x-1\right)&=5\\ \sin3x-1&=5\cos3x \end{align*}\] might be easier to work with.

OpenStudy (jhannybean):

Ohh I see I see.

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ -10\tan(3x) + 25 =\sec^2 (3x)-\tan^2 (3x) \\ -10\tan(3x) +25 =1\\ tan(3x)=\dfrac{12}{5}\\ x=\dfrac{1}{3}arctan\dfrac{12}{5}\\ \approx 0.39}\end{align}\)

OpenStudy (jhannybean):

Oh is that what you were referring to -_- I goofed.

OpenStudy (mathmath333):

@Jhannybean there should be\(+25\) in ur equation

OpenStudy (jhannybean):

Yes you are right :) Haha.

ganeshie8 (ganeshie8):

\[\rm \tan(3x)=\dfrac{12}{5} \implies 3x = \arctan\left(\frac{12}{5}\right) + n\pi \implies x = \frac{1}{3}\left[ \arctan\left(\frac{12}{5}\right) + n\pi\right]\]

OpenStudy (mathmath333):

why not \(\Large 2n\pi\)

ganeshie8 (ganeshie8):

period of tan is just \(\pi\), so it repeats after every \(\pi\) : \[\tan(x) = \tan(x+\pi) = \tan(x+\pi+\pi) = \tan(x+\pi+\pi+\pi+\cdots)\]